A 1.9 m radius cylinder with a mass of 531.1 kg rolls without slipping down a hill which is 56.5 meters high. At the bottom of the hill, what fraction of its total kinetic energy is invested in rotational kinetic energy?
Answer #1

si 9-MSAT sy 212238 = 91.3 mjöx (186*(1-0) T=555*c final temparature. 8= 1.am m = 531.1 kg h 56.5m @VP? 1 = 4 mm 5,4x531-4x0

Request: Dear sir/madam, Please rate it. It will help to improve my CF score. Please thanks in advance.